prove with inductions that n^2+n divisible by 2


1
Expert's answer
2020-12-16T19:06:15-0500

To show that "n^2+n" is divisible by 2, we must show that "n^2+n=2M" for some "M \in \mathbb{Z}" by induction.


For "n=1", we have "1^2+1=2(1)." Here, "M=1\in \mathbb{Z}" . So, it is true for "n=1" .

Assume it is true for "n=k, k\in \mathbb{Z}^+" . That is,


"k^2+k=2M .....................................(1)"


For "n=k+1" . We have;


"(k+1)^2+(k+1)=k^2+2k+1+k+1\\
=k^2+3k+2....................................(2)"


From (1), "k^2=2M-k." Put this into (2)


"=2M-k+3k+2\\
=2M+2k+2\\
=2(M+k+1)\\
=2M'" where "M'=(M+k+1)\in\mathbb{Z}" .

Therefore, it is true for "n=k+1" .


Hence, "n^2+n" is divisible by 2.